# Randomized Maximum Likelihood via High-Dimensional Bayesian Optimization

### v2

@@ 2. Sample $\mu_{n}\sim{\mathcal{N}}_{D}(\mu,\Sigma)$ from the Gaussian prior
Here, we use the notation $[n_{RML}]=\{1,2,\ldots,n_{RML}\}$. Thus, RML solves $n_{RML}$ optimization problems, each with a different objective function, $O_{n}(x)$. The resulting solutions $\{x^{\star}_{n}\}_{n=1}^{n_{RML}}$ are approximate samples from the posterior distribution $p(x|\mathcal{D})$; the samples are only exact draws from the posterior when the simulator $f(x)$ is linear [5]. Nonetheless, good practical performance has been observed for nonlinear (deep) neural network parametrized simulators [6]; see also [7] for a modification of RML for multi-modal posteriors with highly nonlinear simulators.
@@ 3.1 Results
Refer to caption Elliptic PDE
Refer to caption Ebola
Refer to caption MHD
Refer to caption HIV
@@
Refer to caption PDE
Refer to caption RML oracle
Refer to caption HDBO-RML
Refer to caption Ebola
Refer to caption RML oracle
Refer to caption HDBO-RML
Refer to caption MHD
Refer to caption RML oracle
Refer to caption HDBO-RML
Refer to caption HIV
Refer to caption RML oracle
Refer to caption HDBO-RML

### v3

@@ 2. Sample $\mu_{n}\sim{\mathcal{N}}_{D}(\mu,\Sigma)$ from the Gaussian prior
Here, we use the notation $[n_{RML}]=\{1,2,\ldots,n_{RML}\}$. Thus, RML solves $n_{RML}$ optimization problems, each with a different objective function, $O_{n}(x)$. The resulting solutions $\{x^{\star}_{n}\}_{n=1}^{n_{RML}}$ are approximate samples from the posterior distribution $p(x|\mathcal{D})$. The samples are exact draws from the posterior if the simulator $f(x)$ is linear; see Appendix A for a simple proof, or [6] for the original proof treating a more general case. Nonetheless, good practical performance has been observed for nonlinear (deep) neural network parametrized simulators [7]; see also [8] for a modification of RML for multi-modal posteriors with highly nonlinear simulators.
@@ 3.1 Results
Elliptic PDE
Ebola
MHD
HIV
@@
PDE
RML oracle
HDBO-RML
Ebola
RML oracle
HDBO-RML
MHD
RML oracle
HDBO-RML
HIV
RML oracle
HDBO-RML
@@ 4 Discussion
Appendix A Proof for the linear simulator case
In case of a linear simulator $f(x):=Hx$, Gaussian prior distribution $x\sim{\mathcal{N}}(\mu,\Sigma)$, and Gaussian likelihood $\mathcal{D}|x\sim{\mathcal{N}}(Hx,\Sigma_{\text{obs}})$, the posterior distribution $p(x|\mathcal{D})$ is also Gaussian:
$x|\mathcal{D}\sim{\mathcal{N}}(m,V),$
where
$V$ $=(\Sigma^{-1}+H^{\top}\Sigma_{\text{obs}}^{-1}H)^{-1}$
[…]
To see why Randomized Maximum Likelihood (RML) from Algorithm 1 produces exact samples from the posterior distribution in this case, we first recall the RML objective functions from (1):
$O_{n}(x):=\log{\mathcal{N}}_{m}(Hx|\mathcal{D}_{n},\Sigma_{\text{obs}})+\log{\mathcal{N}}_{D}(x|\mu_{n},\Sigma),$
where $\mu_{n}:=\mu+\epsilon_{n}$ and $\mathcal{D}_{n}:=\mathcal{D}+\eta_{n}$, with $\epsilon_{n}\sim{\mathcal{N}}(0,\Sigma)$ and $\eta_{n}\sim{\mathcal{N}}(0,\Sigma_{\text{obs}})$ independently.
Differentiating with respect to $x$ gives
$\nabla O_{n}(x)=2\Sigma^{-1}x-2\Sigma^{-1}\mu_{n}+2H^{\top}\Sigma_{\text{obs}}^{-1}Hx-2H^{\top}\Sigma_{\text{obs}}^{-1}\mathcal{D}_{n}$
and setting this equal to zero and rearranging gives
$x$ $=(\Sigma^{-1}+H^{\top}\Sigma_{\text{obs}}^{-1}H)^{-1}(\Sigma^{-1}\mu_{n}+H^{\top}\Sigma_{\text{obs}}^{-1}\mathcal{D}_{n})$
$=m+V(\Sigma^{-1}\epsilon_{n}+H^{\top}\Sigma_{\text{obs}}^{-1}\eta_{n}).$
The distribution of $x$ can then easily be seen to be
$x$ $\sim{\mathcal{N}}(m,V(\Sigma^{-1}\Sigma\Sigma^{-1}+H^{\top}\Sigma_{\text{obs}}^{-1}\Sigma_{\text{obs}}\Sigma_{\text{obs}}^{-1}H)V^{\top})$
$={\mathcal{N}}(m,V),$
i.e., $x$ is a sample from the true posterior distribution.